Two balls are drawn in succession without replacement from an urn containing 4 red balls and 5 blue balls. Let R be the random variable representing the number of red balls. Find the values of the random variable R

1
Expert's answer
2022-03-07T08:25:06-0500

Let's denote R - red ball, B - blue ball.

Sample space S is all possible outcomes.


"S=\{BB, BR, RB, RR\}"

The possible values of the random variable "R" are "0, 1, 2."

We will assume that the probability of getting heads and tails is the same:


"p = q =1\/2""\def\arraystretch{1.5}
 \begin{array}{c:c}
 Possible \ Outcomes & R \\ \hline
 BB & 0 \\
 \hdashline
 BR & 1 \\
 \hdashline
 RB & 1 \\
 \hdashline
 RR & 2 \\
\end{array}"


Construct the probability distribution of the random variable


"P(BB)=\dfrac{5}{9}(\dfrac{4}{8})=\dfrac{5}{18}"

"P(BR)=\dfrac{5}{9}(\dfrac{4}{8})=\dfrac{5}{18}"

"P(RB)=\dfrac{4}{9}(\dfrac{5}{8})=\dfrac{5}{18}"

"P(RR)=\dfrac{4}{9}(\dfrac{3}{8})=\dfrac{3}{18}"

"\def\arraystretch{1.5}
 \begin{array}{c:c}
 r & 0 & 1 & 2 \\ \hline
 p(r) & 5/18 & 5/9 & 3/18
\end{array}"